# Atlas-free Brain Network Transformer

### v1

@@ 4.4 Discussion
The proposed atlas-free BNT leverages individualized brain parcellation to mitigate ROI misalignment and enhance functional homogeneity. Functional homogeneity is quantified by computing intra-correlations between voxel-wise BOLD time series and the corresponding mean time series within each ROI. Fig. 7 presents the distributions of average intra-correlations across 100 subjects obtained using different parcellation methods. In atlas-based approaches, a fixed atlas is applied uniformly to all subjects, leading to reduced homogeneity due to inter-subject variability. By contrast, individualized parcellation adapts to each subject’s unique brain organization and allows homogeneity to be tuned by setting parameters such as the merging correlation threshold $\nu$ in agglomerative clustering or the correlation cutoff threshold $\tau$ in spectral clustering, resulting in consistently higher functional homogeneity.
@@
(b) Brain connectome age prediction
Figure 8: Averaged Grad-CAM saliency maps overlayed on the MNI brain template highlight the regional contributions to the atlas-free BNT model’s inference. Warmer colors indicate regions with stronger positive influence on the output, revealing the brain areas most relevant to sex classification and brain connectome age prediction.
@@
The Grad-CAM map for the brain-connectome age model in Fig. 8(b) shows widespread, high saliency across the cortical gray-matter mantle—spanning frontal, parietal, temporal, and occipital cortices, including the medial wall—and equally strong saliency in subcortical nuclei (thalamus/striatal complexes) and the posterior cerebellum (vermis and hemispheric lobules). This pan-cortical plus subcortico-cerebellar pattern indicates that age estimates are derived from a distributed cortico–subcortical–cerebellar network rather than a focal locus, consistent with broad lifespan effects including global cortical thinning [35], frontostriatal and thalamo-cortical connectivity reconfiguration with aging [36, 37, 38], and cerebellar structural change across adulthood [39]. We emphasize that Grad-CAM highlights supportive evidence within this trained model (not causal mechanisms), and that block-wise reconstruction and cross-subject averaging may smooth boundaries; nevertheless, the convergence of strong cortical, subcortical, and cerebellar saliency supports a network-level basis for the model’s age predictions.
@@ 5 Conclusion
Future work may further enhance our method by integrating multimodal neuroimaging data, such as structural MRI or diffusion MRI, to provide a more comprehensive characterization of individual brain networks. Additionally, adapting the framework to longitudinal data could enable the modeling of dynamic functional connectivity patterns over time, potentially improving the sensitivity of clinical biomarkers for neurological and psychiatric disorders. Finally, applying and validating our approach across diverse clinical populations could further establish its utility and robustness in personalized medicine applications.
@@ Appendix A Parameter Tuning
The elastic net involves tuning two regularization hyperparameters, $\lambda_{1}$ and $\lambda_{2}$ for the $l_{1}$-norm and $l_{2}$-norm regularizations respectively. We set $\lambda_{1}=1e^{-5}$, $\lambda_{2}=5e^{-5}$ in sex classification, and $\lambda_{1}=0.04$, $\lambda_{2}=0$ in brain connectome age prediction. For the BrainGNN, we conducted a grid search over learning rate $\{0.001,0.005,0.01\}$, feature dimension $\{100,200,300,500\}$, and number of ROIs $\{50,100,200,300\}$; the best validation setting was learning rate 0.005, feature dimension 200, and 200 ROIs. For Graphormer, we set the learning rate to $1e^{-4}$ and used two encoder layers, four attention heads, and 256-dimensional embeddings. For BNT, we used the default learning rate scheduling strategy, two encoder layers, four attention heads, and 360-dimensional embeddings.
@@ Appendix B Brain Connectome Age Prediction
Figure 9: The predicted brain connectome age versus chronological age by the atlas-based approaches with their best performing atlases: a) Elastic net (Craddock-400), b) BrainGNN (AAL), c) Graphormer (Craddock-400), d) BNT (Craddock-400). Each point represents one subject/session. The mean and standard deviation of the absolute errors are calculated for all methods: a) $4.86\pm 3.54$ years; b) $4.63\pm 3.47$ years; c) $4.71\pm 3.64$; d) $4.21\pm 3.21$.

### v2

@@ 4.5 Discussion
The proposed atlas-free BNT leverages individualized brain parcellation to mitigate ROI misalignment and enhance functional homogeneity. Functional homogeneity is quantified by computing intra-correlations between voxel-wise BOLD time series and the corresponding mean time series within each ROI. Fig. 7 presents the distributions of average intra-correlations across 100 subjects obtained using different parcellation methods. In atlas-based approaches, a fixed atlas is applied uniformly to all subjects, leading to reduced homogeneity due to inter-subject variability. By contrast, individualized parcellation adapts to each subject’s unique brain organization and allows homogeneity to be tuned by setting parameters such as the merging correlation threshold $\nu$ in agglomerative clustering (AC) or the correlation cutoff threshold $\tau$ and the number of clusters $k$ in spectral clustering (SC), resulting in consistently higher functional homogeneity. Notably, SC homogeneity can be further increased by using a larger $k$, which typically yields smaller ROIs with more homogeneous signals; in our experiments, we set $k=400$ to achieve the best overall predictive performance.
AC and SC produce individualized parcellations through fundamentally different mechanisms, which leads to distinct ROI configurations. AC is effectively deterministic given a fixed similarity measure and stopping rule, since the greedily merged sequence is uniquely determined. By comparison, SC introduces randomness through the $k$-means step used to cluster the spectral embeddings; different initializations can lead to different solutions, and occasional poor local optima may require rerunning $k$-means with a new initialization to obtain stable parcellations.
@@
(b) Brain-connectome age prediction
Figure 8: Averaged Grad-CAM saliency maps overlayed on the MNI brain template highlight the regional contributions to the atlas-free BNT model’s inference. Warmer colors indicate regions with stronger positive influence on the output, revealing the brain areas most relevant to sex classification and brain-connectome age prediction.
@@
The Grad-CAM map for the brain-connectome age model in Fig. 8(b) shows widespread, high saliency across the cortical gray-matter mantle (spanning frontal, parietal, temporal, and occipital cortices, including the medial wall) and equally strong saliency in subcortical nuclei (thalamus/striatal complexes) and the posterior cerebellum (vermis and hemispheric lobules). This pan-cortical plus subcortico-cerebellar pattern indicates that age estimates are derived from a distributed cortico–subcortical–cerebellar network rather than a focal locus, consistent with broad lifespan effects including global cortical thinning [8], frontostriatal and thalamo-cortical connectivity reconfiguration with aging [35, 5, 9], and cerebellar structural change across adulthood [1]. We emphasize that Grad-CAM highlights supportive evidence within this trained model (not causal mechanisms), and that block-wise reconstruction and cross-subject averaging may smooth boundaries; nevertheless, the convergence of strong cortical, subcortical, and cerebellar saliency supports a network-level basis for the model’s age predictions. It is also reassuring to observe that aggolerative and spectral clustering approaches provide similar salience maps for both sex and age prediction, suggesting these results are more likely to be from intrinsic biological effects rather than artifacts of the specific technical implementations employed.
@@ 5 Conclusion
We also acknowledge several limitations of the proposed atlas-free BNT and plan to address them in future work. Although the method is designed to be robust to imperfect alignment, it still relies on spatial normalization to a common template in the MNI space; we will therefore investigate strategies that further reduce sensitivity to misregistration, including improved alignment quality control and more deformation-tolerant feature aggregation. Because individualized parcellation depends on clustering hyperparameters (e.g., $\nu$, $\tau$ and $k$), we will study principled selection criteria and stability analyses to reduce sensitivity to these choices and improve reproducibility across datasets/tasks. We also plan to optimize the computational pipeline to facilitate deployment on larger cohorts. Finally, we will conduct broader multi-site and clinical validations, and explore domain-adaptation strategies, to ensure that atlas-free BNT generalizes reliably across acquisition settings and diverse patient populations.
6 Acknowledgement
The study was supported by the National Institutes of Health grants/awards (R01AG089806, R01AG070937, R01AG072603, and P30AG066511).
@@ Appendix A Parameter Tuning
To ensure a fair comparison against the four atlas-based baselines, we tuned their hyperparameters following the procedures summarized below:
  • For the elastic net, two regularization hyperparameters, $\lambda_{1}$ and $\lambda_{2}$, control the strengths of the $l_{1}$- and $l_{2}$-norm penalties, respectively. We set $\lambda_{1}=1e^{-5}$ and $\lambda_{2}=5e^{-5}$ for the sex-classification task, and $\lambda_{1}=0.04$ and $\lambda_{2}=0$ for brain-connectome age prediction.
  • For BrainGNN, we performed a grid search over learning rate $\{0.001,0.005,0.01\}$, feature dimension $\{100,200,300,500\}$, and number of ROIs $\{50,100,200,300\}$; the best validation configuration was learning rate $0.005$, feature dimension $200$, and $200$ ROIs.
  • For Graphormer, we set the learning rate to $1e^{-4}$ and used two encoder layers, four attention heads, and 256-dimensional embeddings.
  • For the original BNT, we adopted the default cosine-annealing learning-rate scheduling strategy that starts with $1e^{-4}$, and used two encoder layers, four attention heads, and 360-dimensional embeddings.
@@ Appendix B Brain Connectome Age Prediction
Fig. 9 shows the four atlas-based approaches’ predicted brain-connectome age plotted against chronological age.
@@
Figure 9: The predicted brain-connectome age versus chronological age by the atlas-based approaches with their best performing atlases: a) Elastic net (Craddock-400), b) BrainGNN (AAL), c) Graphormer (Craddock-400), d) BNT (Craddock-400). Each point represents one subject/session. The mean and standard deviation of the absolute errors are calculated for all methods: a) $4.86\pm 3.54$ years; b) $4.63\pm 3.47$ years; c) $4.71\pm 3.64$; d) $4.21\pm 3.21$.